# RapidPoseTriangulation: Multi-view Multi-person Whole-body Human Pose Triangulation in a Millisecond

### v3

@@ 6 Conclusion
In this work a new algorithm for 3D human pose triangulation under the name of RapidPoseTriangulation was presented. It is based on the idea of triangulating pairs of 2D poses from different views separately and grouping and merging them to the final poses directly in 3D space. While the general approach is relatively simple, it is also much faster than any previous algorithm. Unlike most other methods, it is also able to predict whole-body poses.
@@
In a broad evaluation of the algorithm on several datasets it showed state-of-the-art generalization performance. It scales well with an increased number of cameras and is especially suited for real-time applications. Besides that, the results also show, that a relatively simple algebraic approach still can outperform most modern fully differentiable methods in the task of exploiting multi-view information.
@@
In summary, it is one of the best and by far the fastest method for multi-view multi-person human pose estimation that is currently available. By making the source-code publicly available, it is hoped that this work can contribute to the development of more intuitive and safer human-computer or human-robot interactions.

### v4

@@ 6 Conclusion
This work presented RapidPoseTriangulation, a new algorithm for 3D human pose triangulation. Its basic idea is to triangulate pairs of 2D poses from different views separately, and group and merge them to the final poses directly in 3D space. This relatively simple concept makes it very efficient, and much faster than any previous algorithm.
@@
In a broad evaluation of the algorithm on several datasets it showed state-of-the-art performance. It scales well with an increased number of cameras and is especially suited for real-time applications. Besides that, the results also show that a relatively simple algebraic approach still can outperform many recent fully differentiable or end-to-end learned multi-view methods. This suggests that the current trend of increasing model complexity does not necessarily translate into better exploitation of the geometric structure.
@@
In summary, it is one of the best and by far the fastest method for multi-view multi-person human pose estimation that is currently available. In contrast to most prior works, it is also able to predict whole-body poses, which can be used to implement more advanced follow-up tasks. By making the source-code publicly available, it is hoped that this work can contribute to the development of more intuitive and safer human-computer or human-robot interactions.
Appendix A Ablation Studies
The following section contains the ablation studies that were left out of the main paper for space reasons.
A.1 Number of cameras
For completeness an experiment with panoptic using , , , , or cameras, was executed. As expected, the results show, that more cameras lead to better results. The frame-rate also slows down, with the 2D-part having the most impact, and the 3D-part having a slight exponential increase, even though its impact is still very small.
For real-world applications this experiment is not very meaningful though, since the required number of cameras is highly dependent on the targeted environment. Depending on the setup, cameras can be enough or much too few, as in campus versus mvor. VP and F-VP both did an experiment for cameras on panoptic in which they had only very small performance drops, but their performance on campus and mvor showed a large drop, even though they also have cameras. For a real-world application, it therefore is better to first select a similar dataset, and then evaluate the impact of the number of cameras, and also their best positioning. Since it’s not possible to evaluate all combinations of datasets, cameras and positions, the recommendation for developers is instead to use the provided source-code and run a personalized evaluation on the most similar dataset.
| Number of cameras | PCP | PCK@100/500 | MPJPE | Recall@100/500 | Invalid | F1 | Time-3D | FPS | ||
| 3 | 90.4 | 83.7 | 93.5 | 58.2 | 85.2 | 94.5 | 1.1 | 96.7 | 0.1 | 95.8 |
| 5 | 99.1 | 96.5 | 99.8 | 30.6 | 99.3 | 99.8 | 1.8 | 99.0 | 0.2 | 57.5 |
| 7 | 99.5 | 97.2 | 99.9 | 27.9 | 99.9 | 99.9 | 3.7 | 98.1 | 0.3 | 40.9 |
| 10 | 99.7 | 98.0 | 99.9 | 23.9 | 99.9 | 99.9 | 3.0 | 98.4 | 0.5 | 29.6 |
| 31 | 99.8 | 98.7 | 99.9 | 19.3 | 99.9 | 99.9 | 2.8 | 98.6 | 5.0 | 9.1 |
| Ablation | Dataset | PCP | PCK@100/500 | MPJPE | Recall@100/500 | Invalid | F1 | Time-3D | ||
| default | shelf | 99.2 | 94.4 | 100 | 47.5 | 98.7 | 100 | 41.3 | 74.0 | 0.1 |
| score without 2D confidence | shelf | 99.2 | 94.4 | 100 | 47.7 | 98.7 | 100 | 41.3 | 74.0 | 0.1 |
| no pair pre-filtering | shelf | 99.2 | 94.4 | 100 | 47.6 | 98.7 | 100 | 40.4 | 74.6 | 0.2 |
| keep topk outliers | shelf | 98.9 | 93.8 | 100 | 49.8 | 97.9 | 100 | 41.3 | 74.0 | 0.1 |
| keep topk+distance outliers | shelf | 98.9 | 93.7 | 100 | 49.9 | 97.9 | 100 | 41.3 | 74.0 | 0.1 |
| min_match_score=0.8 | shelf | 99.2 | 94.5 | 100 | 47.7 | 97.9 | 100 | 46.9 | 69.3 | 0.1 |
| default | panoptic | 99.1 | 96.5 | 99.8 | 30.6 | 99.3 | 99.8 | 1.8 | 99.0 | 0.2 |
| no postprocessing steps | panoptic | 99.0 | 96.3 | 99.8 | 31.6 | 99.2 | 99.8 | 1.8 | 99.0 | 0.1 |
| default | badminton | 99.9 | 99.5 | 100 | 24.3 | 100 | 100 | 0.0 | 100 | 0.7 |
| only triangulation error | badminton | 99.9 | 99.6 | 100 | 24.6 | 100 | 100 | 0.0 | 100 | 0.5 |
| only reprojection error | badminton | 100 | 99.5 | 100 | 24.4 | 100 | 100 | 0.0 | 100 | 0.7 |
| min_group_size=1 | badminton | 99.4 | 98.9 | 99.8 | 27.1 | 99.3 | 100 | 39.0 | 75.8 | 1.0 |
| default | tagging | 97.5 | 92.9 | 99.4 | 47.9 | 90.0 | 100 | 4.1 | 97.9 | 0.3 |
| no tracking | tagging | 97.5 | 93.0 | 99.3 | 46.7 | 91.8 | 99.8 | 3.9 | 97.9 | 0.3 |
A.2 Algorithmic ablations
In general the algorithm is very robust against small changes, and only a few of them have a notable impact on the results. Ignoring the 2D confidence scores instead to integrating them into the 3D keypoint score, results in a slightly higher MPJPE. Removing the pair pre-filtering step, which removes pairs of poses before the triangulation, has minimal impact on most scores, but notably increases the triangulation time, since more pairs have to be calculated. The two outlier removal steps, before merging the proposals of one group to a single person notably improves the accuracy. Lowering the error threshold, that filters poor triangulations before merging, slightly increases MPJPE and the invalid count. The postprocessing steps of dropping clearly invalid persons or replacing clearly wrong joints by with their neighbors also improves the MPJPE and reduces the invalid person count in some datasets. Using only the triangulation error or only the reprojection error for the 3D keypoint score calculation, instead of a combination of both, has a slightly negative impact on the MPJPE. Using only the triangulation score is slightly faster though, because the more complex reprojection can be skipped. Especially in datasets with more cameras it is recommended to use increase the minimal required group size. This results in dropping of invalid 3D proposals which do not originate from the same person, but have a similar pose in two or more views. The higher the number of cameras the higher the chance for it. Since they are from different persons, they are triangulated to some mathematically valid positions somewhere in the 3D space. Proposals from the same person on the other hand start to build clusters around their real position, by which they can be recognized as such. The tracking step has only a minimal impact on most datasets, since it does not smooth the movement trajectories. At fast movements it might slightly reduce the position accuracy because it can slow down some position changes. On the other hand it can prevent physically impossible fast jitter movements. In case a person is not visible for a few frames, it can also prevent the person from being lost completely, since their last position is returned then, but it will also keep persons moving out of the room alive for a few extra frames. Therefore tracking is considered as an optional extension.